OpenStudy (minatsukisaya95):

how can i solve this..the sum of 2 non-negative number is to be 1. how should these numbers can be chosen so that the sum of their squares is a maximum? minimum?

OpenStudy (yttrium):

I think you should use second derivative test.

OpenStudy (minatsukisaya95):

i have used that but i only got the local min. how to get its local max

OpenStudy (debbieg):

If there is no local max, then you have to check the function value at each endpoint. You have a function in one variable, let's say a, where a is one of the two numbers. Then since a+b=1 and both are non-negative, you have \(0\le a\le 1\). So since your quadratic does not have a max, and only one min, you just need to find it's value for a=1 and a=0.

OpenStudy (minatsukisaya95):

ok thank u it really helps me